OpenStudy (anonymous):

Graph the following equation by plotting three points. -2y=-x+8

OpenStudy (fanduekisses):

first convert to slope intercept form

OpenStudy (anonymous):

I don't know how to do that....

OpenStudy (anonymous):

y=mx+b, where m is the slope and b is the y-intercept. So, solve for y.

OpenStudy (anonymous):

I don't understand how to do that equation tho.

OpenStudy (anonymous):

I'm trying so hard to understand this stuff. I have been working on it for a week already.

OpenStudy (anonymous):

Do you know how to solve for y? Fanduekisses drawing explains it. -2y=-x+8 divide both sides by -2 to get y by itself. -2y=-x+8 First, divide both sides by -2 to get y by itself: \[\frac{ -2y }{ -2 }=\frac{ -x }{ -2 }+\frac{ 8 }{ -2 }\] So you get: \[y=\frac{ 1 }{2 }x-4\] m=1/2, that's the slope of your line -4 is the y-intercept (the place where x is 0 and your line crosses the y axis)